# Price Experimentation and Interference

### v2

@@
In this paper, we examine biases arising in A/B tests where firms modify a continuous parameter, such as price, to estimate the global treatment effect of a given performance metric, such as profit. These biases emerge in canonical experimental estimators due to interference among market participants. We employ structural modeling and differential calculus to derive intuitive characterizations of these biases. We then specialize our general model to a standard revenue management pricing problem. This setting highlights a key pitfall in the use of A/B pricing experiments to guide profit maximization: notably, the canonical estimator for the expected change in profits can have the wrong sign. In other words, following the guidance of canonical estimators may lead firms to move prices in the wrong direction, inadvertently decreasing profits relative to the status quo. We apply these results to a two-sided market model and show how this “change of sign” regime depends on model parameters such as market imbalance, as well as the price markup.Finally, we discuss structural and practical implications for platform operators.
@@ 1 Introduction
We apply the results of the revenue management setting to the two-sided market model of Johari et al., (2022). We show how the biases of listing and customer-side experiments depend on model parameters with a particular focus on regimes in which this “change of sign” occurs. For example, for a listing-side experiment a platform is more likely to obtain the incorrect sign in an experiment if the demand is relatively low. We also calibrate this stylized model to experimental estimates from Airbnb (published in Holtz et al., 2020), and investigate empirical implications for regimes where a change of sign may occur in a price experiment.
@@ 2 Related Work
The phenomenon we study is an interference effect capturing the “cannibalization” of treatment purchases by their control counterparts. More generally, this interference effect is a violation of the Stable Unit Treatment Value Assumption (SUTVA) (Imbens and Rubin, 2015). Interference among users has been widely studied in social networks (Manski, 2013, Athey et al., 2018, Basse et al., 2019, Saveski et al., 2017, see, e.g.) and more recently on two-sided platforms (Blake and Coey, 2014b , Fradkin, 2019, Liu et al., 2020), showing its importance in practice. Platforms have developed alternative designs and/or estimators to alleviate these biases such as cluster randomization (Ugander et al., 2013, Chamandy, 2016, Holtz et al., 2020), switchback testing (Chamandy, 2016, Bojinov et al., 2022, Glynn et al., 2020, Hu and Wager, 2022), two-sided randomization and bipartite experiments (Bajari et al., 2021, Johari et al., 2022, Pouget-Abadie et al., 2019), shadow-price-based methods (Bright et al., 2023), Q-learning (Farias et al., 2022, farias2023correcting), and message passing (shirani2023causal).
A number of other papers prior to ours have considered price experiments specifically, including Wager and Xu, (2021), Munro et al., (2021), Li et al., (2023), Aparicio et al., (2023), roemheld2023interference, Deng and Le, (2023). Both Wager and Xu, (2021) and Li et al., (2023) consider estimation of treatment effects from marginal price changes; in Wager and Xu, (2021), this is studied via a mean-field limit approach that admits a debiasing technique, while in Li et al., (2023), the emphasis is on the estimation of such effects via model-free and model-based estimators in switchback experiments with stochastic congestion (i.e., due to queueing effects). roemheld2023interference uses an empirical study to suggest false positives may be inflated due to interference in pricing experiments. More similar to our findings, Deng and Le, (2023) also notes the possibility that the sign of the treatment effect can flip; in their setting this is due to an effect they call lead-day bias in markets with expiring goods. In an auction design setting, Chawla et al., (2016) points out that A/B testing different auction formats can also lead to the wrong sign. Finally, more broadly, we note that Carneiro et al., (2010) refers to the kinds of treatment effects we consider in this paper – those where a continuous variable (such as price) is incrementally varied – as marginal policy changes, and develops a framework to study them.
@@ 6.2 Steady-State and Estimand
First, we note that listing availability $s^{*}(\lambda,p)$ is monotonically increasing with price $p$ and decreasing with the market balance parameter $\lambda$, which is natural. These monotonicity results can be derived by differentiating $s^{*}(\lambda,p)$ in equation (13). [^7] Indeed, we obtain $s^{*}_{p}(\lambda,p)=-\frac{s^{*}(\lambda,p)v^{\prime}(p)\varepsilon}{\frac{1}{\lambda}(\varepsilon+s^{*}(\beta,p)v(p))^{2}+v(p)\varepsilon}>0$ and $s^{*}_{\lambda}(\lambda,p)=-\frac{v(p)s(\lambda,p)(v(p)s(\lambda,p)+\varepsilon)}{\varepsilon v(p)(2s(\lambda,p)+\lambda)+v(p)^{2}s(\lambda,p)^{2}+\varepsilon^{2}}<0$. Figure 4 in the Appendix illustrates the steady-state listing availability as a function of price ($p$) and market balance ($\lambda$).
@@ 6.3.2 Customer-Side Randomization
$\rho-s^{*}(\lambda,p_{0},p_{1},q)=\lambda\left[q\frac{s^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}{\varepsilon+s^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}+(1-q)\frac{s^{*}(\lambda,p_{0},p_{1},q)v(p_{0})}{\varepsilon+s^{*}(\lambda,p_{0},p_{1},q)v(p_{0})}\right].$ (18)
@@ 6.5 Change of Sign
Theorems 6.3 and 6.4 tell us that regardless of the market balance parameter $\lambda$, there is always an interval of prices $P_{\lambda}$ such that the sign of the canonical estimator will be wrong, regardless of whether a listing-side or customer-side experiment is run. The proofs are provided in Appendix 10.9 and repeatedly exploit the market balance equations for CR and LR respectively.
@@ 6.5.1 Numerical Results
Consistent with Propositions 6.1 and 6.2, we observe that the change of sign region becomes empty in the demand-constrained regime for a CR experiment and in the supply constrained experiment for an LR experiment. At intermediate market balance regimes, we observe that the change of sign region corresponds to a single continuous interval on prices, as the theorems above suggest. We observe that in the demand-constrained regime Condition (b) for a change of sign is more likely to be satisfied by the LR experiments, while in the supply-constrained regime Condition (b) is more likely to be satisfied by the CR experiments as expected. We also see that the region where a change of sign occurs is, overall, increasing with market balance for the CR experiments and decreasing for the LR experiment.
7 An Empirical Case Study
Our results show that price experiments can get the sign wrong, and suggest that practitioners would benefit from investigating whether their experiments are vulnerable to such changes of sign. In this section, we present an empirical illustration to study this question, by calibrating a version of our stylized model to data from Airbnb. Although our model simplifies the complexity of the Airbnb marketplace in many dimensions, we believe this exercise is valuable because it offers practical insights into how our results can align with empirical data. As we will see, our results suggest that a change of sign is indeed plausible in the Airbnb context.
We rely on the estimates by Holtz et al., (2020) of a guest fee “meta-experiment” at Airbnb. This experiment involved a treatment that raised platform fees. The design consisted of two “arms”: randomization across clusters of listings in one arm; as well as Bernoulli randomization across listings in a second arm. Holtz et al., (2020) use this meta-experiment to estimate the degree to which Bernoulli randomized experiments exhibit interference, by comparing with estimates from the cluster randomized arm (which mitigates interference effects). The paper reports on cluster and naive experiment estimates and standard errors for the global treatment effect (the TATE, or total average treatment effect in their paper) for bookings, nights booked, and gross spend.
To translate from Holtz et al., (2020) to our setting, we first recap certain features of their presentation. First, even though the treatment corresponds to a percentage increase in prices, the size of this treatment intervention is not disclosed in Holtz et al., (2020). For our illustrative exercise here, we interpret the estimates of Holtz et al., (2020) as semi-elasticities, i.e., absolute changes in quantities for a 1% price increase. Second, Holtz et al., (2020) does not disclose absolute levels (e.g., of nights booked, or revenue); instead, these quantities are rescaled by the same randomly chosen quantity (not disclosed to the reader). As we will see, this does not present any challenges for our exercise here.
We need to slightly adjust the analysis to accommodate a percentage change in prices relative to the bias analysis in Section 5 that considers an absolute change. Such an analysis can be carried out by considering demand $D(\alpha p)$ and profit $(\alpha p-c)D(\alpha p)$, and considering derivatives with respect to $\alpha$, evaluated at $\alpha=1$. Proceeding in this way, we redefine:
$GTE_{D}(p)=pD^{\prime}(p);\ \ GTE_{\pi}(p)=p(D(p)+pD^{\prime}(p))-cpD^{\prime}(p)$
as the semi-elasticities of demand and profit, respectively.
As before, the first condition for a change of sign at $p_{0}$ is (a) $GTE_{\pi}(p_{0})\geq 0$, which becomes:
$p_{0}(D(p_{0})+p_{0}D^{\prime}(p_{0}))-cp_{0}D^{\prime}(p_{0}).$ (21)
The second condition is that (b) $Bias_{\pi}(p_{0},q)\geq GTE_{\pi}(p_{0})$ (where now we note that the bias is computed using percentage price changes, i.e., with respect to changes in $\alpha$). Since $Bias_{\pi}(p_{0},q)=(p_{0}-c)Bias_{D}(p_{0},q)$, condition (b) becomes equivalent to:
$(p_{0}-c)Bias_{D}(p_{0},q)\geq p_{0}(D(p_{0})+p_{0}D^{\prime}(p_{0}))p_{0}-cp_{0}D^{\prime}(p_{0}).$ (22)
We now leverage the empirical estimates of Holtz et al., (2020). The cluster arm estimates are a decrease of $1.136$ in nights booked per listing, and of $116.67$ in gross revenues. For our analysis, we assume that the cluster arm provides ground truth, and the corresponding bias for demand is reported to be $0.205$. We assume that $D(p_{0})$ represents nights booked per listing and $p_{0}$ the average price per night, so that $p_{0}D(p_{0})$ represents gross revenues per listing. In particular, note that then $p_{0}D^{\prime}(p_{0})$ is the semi-elasticity of nights booked, and $p_{0}(D(p_{0})+p_{0}D^{\prime}(p_{0}))$ is the semi-elasticity of gross revenues per listing. Finally, we assume $c$ represents the marginal (opportunity) cost per night.
By equation (21) and the estimates, the first condition for change of sign amounts to $-116.7+1.136c\geq 0$, because we identify $-116.67$ and $-1.136$ as the ground-truth semi-elasticities of revenues and demand, respectively. This simplifies to
$c\geq 102.73.$ (23)
The second condition for change of sign amounts to $0.205(p_{0}-c)\geq-116.67+1.136c$, which can be simplified to
$116.67+0.205p_{0}\geq 1.341c$ (24)
Importantly, note that all quantities used in these inequalities are re-scaled by the same random constant, and so these inequalities are valid for any value of that constant under our assumptions.
The pair of inequalities (23)-(24) can be used to interrogate whether a change of sign is plausible in this setting. For example, for a reasonable marginal opportunity cost of $c=125$ (corresponding to \$125 per night), our analysis suggests that market segments with a baseline price of $p_{0}=249$ (corresponding to \$249 per night) yield an incorrectly estimated sign. More generally, this exercise allows us to identify combinations of prices and costs for which the canonical estimator would result in a change of sign. In particular, (24) can be rewritten as $0.847-87.002/p_{0}\leq(p_{0}-c)/p_{0}$, lower bounding markups based on observed market prices that yield a change of sign. For example, the average price of recently booked Airbnb listings in San Francisco is given by \$212 (cf http://insideairbnb.com/san-francisco). According to our analysis, this implies a markup of $0.43$ or larger would yield a change of sign. Given that these are markups that are reasonable in practice, we believe this stylized exercise motivates further detailed empirical study of the prevalence of such changes of sign in real-world platforms.

### v3

@@
In this paper, we examine biases arising in A/B tests where firms modify a continuous parameter, such as price, to estimate the global treatment effect of a given performance metric, such as profit. These biases emerge in canonical experimental estimators due to interference among market participants. We employ structural modeling and differential calculus to derive intuitive characterizations of these biases. We then specialize our general model to a standard revenue management pricing problem. This setting highlights a key pitfall in the use of A/B pricing experiments to guide profit maximization: notably, the canonical estimator for the expected change in profits can have the wrong sign. In other words, following the guidance of canonical estimators may lead firms to move prices in the wrong direction, inadvertently decreasing profits relative to the status quo. We introduce a novel debiasing technique for such experiments, that only requires that the firm equally split experimental units between treatment and control. We apply these results to a two-sided market model and show how this “change of sign” regime depends on model parameters such as market imbalance, as well as the price markup. We also extend our revenue management model to an alternative specification in which firms set a fixed sales fee across heterogeneous products, and show that all our results extend to this setting as well.
[…]
@@ 1 Introduction
We take advantage of our structural characterization of bias to construct a novel debiasing method for the canonical A/B experiment. In particular, we show that if the experiment equally allocates units between treatment and control, then it is possible to fully debias the canonical estimator, by taking advantage of symmetry in the structure of bias. Notably, this allows firms to run the most simple version of an A/B pricing experiment, while still emerging with an unbiased estimator (under a set of assumptions that me make clear below). This is a stark contrast to other methods for reducing bias in A/B experiments such as clustered or switchback experiments, which require much more complicated experimental designs and/or estimators (Holtz et al., 2020, Chamandy, 2016, Bojinov et al., 2022, Glynn et al., 2020, Hu and Wager, 2022).
To obtain additional insight, we apply the results of the revenue management setting to the two-sided market model of Johari et al., (2022). We show how the biases of listing and customer-side experiments depend on model parameters with a particular focus on regimes in which this “change of sign” occurs. For example, for a listing-side experiment, a platform is more likely to obtain the incorrect sign in an experiment if the demand is relatively low. We also investigate numerically how bias depends on both market balance and the baseline control price.
We conclude the paper with an alternative specialization of our revenue management setting to a model where the firm sets a fixed fee as a percentage of overall transaction volume, as is common in online marketplaces. In this model we allow the possibility of a heterogeneous collection of goods for sale, with an unknown (joint) demand function. We show that all our previous results extend to this setting as well.
@@ 2 Related Work
The phenomenon we study is an interference effect capturing the “cannibalization” of treatment purchases by their control counterparts. More generally, this interference effect is a violation of the Stable Unit Treatment Value Assumption (SUTVA) (Imbens and Rubin, 2015). Interference among users has been widely studied in social networks (Manski, 2013, Athey et al., 2018, Basse et al., 2019, Saveski et al., 2017, see, e.g.) and more recently on two-sided platforms (Blake and Coey, 2014b , Fradkin, 2019, Liu et al., 2020), showing its importance in practice. Platforms have developed alternative designs and/or estimators to alleviate these biases such as cluster randomization (Ugander et al., 2013, Chamandy, 2016, Holtz et al., 2020), switchback testing (Chamandy, 2016, Bojinov et al., 2022, Glynn et al., 2020, Hu and Wager, 2022), two-sided randomization and bipartite experiments (Bajari et al., 2021, Johari et al., 2022, Pouget-Abadie et al., 2019), shadow-price-based methods (Bright et al., 2023), Q-learning (Farias et al., 2022, 2023), and message passing (shirani2023causal).
A number of other papers prior to ours have considered price experiments specifically, including Wager and Xu, (2021), Munro et al., (2021), Li et al., (2023), Aparicio et al., (2023), roemheld2023interference, Deng and Le, (2023). Both Wager and Xu, (2021) and Li et al., (2023) consider estimation of treatment effects from marginal price changes; in Wager and Xu, (2021), this is studied via a mean-field limit approach that admits a debiasing technique, while in Li et al., (2023), the emphasis is on the estimation of such effects via model-free and model-based estimators in switchback experiments with stochastic congestion (i.e., due to queueing effects). Roemheld and Rao, (2024) uses an empirical study to suggest false positives may be inflated due to interference in pricing experiments. More similar to our findings, Deng and Le, (2023) also notes the possibility that the sign of the treatment effect can flip; in their setting this is due to an effect they call lead-day bias in markets with expiring goods. In an auction design setting, Chawla et al., (2016) points out that A/B testing different auction formats can also lead to the wrong sign. Finally, more broadly, we note that Carneiro et al., (2010) refers to the kinds of treatment effects we consider in this paper – those where a continuous variable (such as price) is incrementally varied – as marginal policy changes, and develops a framework to study them.
@@ 5 Biases: Analysis and Results
6 Debiasing Canonical Estimators
In this section, we take advantage of our structural characterization of bias to introduce a novel approach to debiasing $\widehat{GTE}_{D}$ and $\widehat{GTE}_{\pi}$ that applies as long as the platform has equally allocated units between treatment and control, i.e., $q=1/2$.
Recall (10) (11) provide the expressions for bias in terms of derivatives of demand. We now show how we can use observables to estimate this bias. First, we show that $D_{0y}(p_{0},p_{0},q)$ can be computed (approximately) using observed data. To see this, note that $D_{0y}(p_{0},p_{0},q)$ is defined by:
$D_{0y}(p_{0},p_{0},q)=\lim_{p_{1}\to p_{0}^{+}}\frac{D_{0}(p_{0},p_{1},q)-D_{0}(p_{0},p_{0},q)}{p_{1}-p_{0}}.$
By Assumption 6 we have that $D_{0}(p_{0},p_{0},q)=D(p_{0})$. Since we assume that the firm was pricing at $p_{0}$ before the experiment began, they have observed $D(p_{0})$ prior to the experiment. The quantity $D_{0}(p_{0},p_{1},q)$ is observed as well: it represents the control demand during the experiment. This means that in the limit as $p_{1}$ approaches $p_{0}$, observables can be used to approximately compute the derivative $D_{0y}(p_{0},p_{0},q)$.
We now note that by symmetry, we have $D_{0}(p_{0},p_{1},1/2)=D_{1}(p_{1},p_{0},1/2)$. This holds because both terms represent demand for the half of the goods that are priced at $p_{0}$, when the other half of goods are priced at $p_{1}$. Taking the derivative of both sides of this equality with respect to $p_{1}$, we obtain $D_{0y}(p_{0},p_{1},1/2)=D_{1x}(p_{1},p_{0},1/2)$, so that when $p_{1}=p_{0}$, there holds $D_{0y}(p_{0},p_{0},1/2)=D_{1x}(p_{0},p_{0},1/2)$. We have shown the following result.
Proposition 6.1
Suppose that $q=1/2$. Then there holds:
$Bias_{D}(p_{0},1/2)$ $=2D_{0y}(p_{0},p_{0},1/2);$ (14)
$Bias_{\pi}(p_{0},1/2)$ $=2(p_{0}-c)D_{0y}(p_{0},p_{0},1/2).$ (15)
In particular, there holds:
$GTE_{D}$ $=\widehat{GTE}_{D}(p_{0},1/2)+2D_{0y}(p_{0},p_{0},1/2);$ (16)
$GTE_{\pi}$ $=\widehat{GTE}_{\pi}(p_{0},1/2)+2(p_{0}-c)D_{0y}(p_{0},p_{0},1/2).$ (17)
Thus if the firm is able to estimate $D_{0y}(p_{0},p_{0},q)$ using observables, then the right hand sides of (16) and (17) are debiased estimators of $GTE_{D}$ and $GTE_{\pi}$ respectively.
We note several limitations of this method. First, of course, the symmetry property being exploited only applies when the market is equally allocated to treatment and control. Second, the assumption that $D(p_{0})$ is observed relies on market stationarity; if the market is nonstationary, this adjusted estimator may not capture the true treatment effect during the experiment. Finally, estimation of $D_{0y}$ uses only half of the experimental data, namely, only the observed control demand, which may be a source of sample inefficiency (i.e., increased variance).
@@ 7.2 Steady-State and Estimand
First, we note that listing availability $s^{*}(\lambda,p)$ is monotonically increasing with price $p$ and decreasing with the market balance parameter $\lambda$, which is natural. These monotonicity results can be derived by differentiating $s^{*}(\lambda,p)$ in equation (18). [^7] Indeed, we obtain $s^{*}_{p}(\lambda,p)=-\frac{s^{*}(\lambda,p)v^{\prime}(p)\varepsilon}{\frac{1}{\lambda}(\varepsilon+s^{*}(\lambda,p)v(p))^{2}+v(p)\varepsilon}>0$ and $s^{*}_{\lambda}(\lambda,p)=-\frac{v(p)s(\lambda,p)(v(p)s(\lambda,p)+\varepsilon)}{\varepsilon v(p)(2s(\lambda,p)+\lambda)+v(p)^{2}s(\lambda,p)^{2}+\varepsilon^{2}}<0$. Figure 4 in the Appendix illustrates the steady-state listing availability as a function of price ($p$) and market balance ($\lambda$).
@@ 7.3.2 Customer-Side Randomization
$\rho-s^{*}(\lambda,p_{0},p_{1},q)=\left[q\frac{s^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}{\varepsilon+s^{*}(\lambda,p_{0},p_{1},q)v(p_{1})}+(1-q)\frac{s^{*}(\lambda,p_{0},p_{1},q)v(p_{0})}{\varepsilon+s^{*}(\lambda,p_{0},p_{1},q)v(p_{0})}\right].$ (23)
@@ 7.5 Change of Sign
Theorems 7.3 and 7.4 tell us that regardless of the market balance parameter $\lambda$, there is always an interval of prices $P_{\lambda}$ such that the sign of the canonical estimator will be wrong, regardless of whether a listing-side or customer-side experiment is run. The proofs are provided in Appendix 11.9 and repeatedly exploit the market balance equations for CR and LR respectively.
@@ 7.5.1 Numerical Results
Consistent with Propositions 7.1 and 7.2, we observe that the change of sign region becomes empty in the demand-constrained regime for a CR experiment and in the supply constrained experiment for an LR experiment. At intermediate market balance regimes, we observe that the change of sign region corresponds to a single continuous interval on prices, as the theorems above suggest. We observe that in the demand-constrained regime Condition (b) for a change of sign is more likely to be satisfied by the LR experiments, while in the supply-constrained regime Condition (b) is more likely to be satisfied by the CR experiments as expected. We also see that the region where a change of sign occurs is, overall, increasing with market balance for the CR experiments and decreasing for the LR experiment.
8 Fee Model
In this section we consider an environment in which a firm charges a fixed sales fee, as opposed to setting a price and incurring a marginal cost for each purchase. This represents the common situation–particularly in two-sided online marketplaces–where firms do not directly control the prices of the goods on their platforms, but instead collect a fixed fee for the sales of heterogeneous goods.
8.1 Specifications
Consider $N$ products, with a fixed vector of prices $\vec{p}\in\mathbb{R}^{N}$. The demand function $D:\mathbb{R}^{N}\to\mathbb{R}^{N}$, takes in a vector of prices and outputs a vector in $\mathbb{R}^{N}$ that represents the quantity demanded for each good. We assume that the platform selects a single fee for all products $f\in\mathbb{R}^{+}$; this fee represents a percentage of all money transacted that the platform will collect. We assume that this fee is passed on entirely to the customer and that the platform’s marginal cost is zero. Therefore, we can write the platform’s profit $\Pi(f,\vec{p})$ as follows: [^10] Prices may be a function of the fee, since firms may pass-through fee changes. However, since experiments tend to be short-lived we ignore these longer term calibration effects.
$\Pi(f,\vec{p})=f\vec{p}\cdot D((1+f)\vec{p}).$ (26)
Let $D^{\prime}(\cdot)$ represent the matrix of demand derivatives in price. We can then write the platform’s profit derivative in the fee as:
$\Pi_{f}(f,\vec{p})=\vec{p}\cdot D((1+f)\vec{p})+f\vec{p}\cdot\vec{p}\cdot D^{\prime}((1+f)\vec{p}).$ (27)
Define $R(f,\vec{p})$ as the total money transacted in the platform when the fee is $f$ and the prices are $\vec{p}$. We can then write $R(f,\vec{p})$ as
$R(f,\vec{p})=\vec{p}\cdot D((1+f)\vec{p}).$
Let $R_{f}(f,\vec{p})$ represent the derivative of this quantity with respect to the fee. We can then rewrite (27) as:
$\Pi_{f}(f,\vec{p})=R(f,\vec{p})+fR_{f}(f,\vec{p}).$ (28)
And rewrite (26) as:
$\Pi(f,\vec{p})=fR(f,\vec{p}).$
{assumption}
We assume that $\Pi_{f}$ exists and is well defined for all $f\in\mathbb{R}^{+}$ and all $\vec{p}\in\mathbb{R}^{N}$, or in other words, that $\Pi$ is differentiable over its domain.
When running an A/B experiment, the platform will compare a control fee $f_{0}$ to a treatment fee $f_{1}$. To keep the presentation succinct, in this section we focus only on estimation of profits. This would cause a true global treatment effect (GTE) of $\Pi(f_{1},\vec{p})-\Pi(f_{0},\vec{p})$. Thus we define the (infinitesimal) $GTE$ at the control fee $f_{0}$ as:
$GTE_{\Pi}(f_{0},\vec{p})=\Pi_{f}(f_{0},\vec{p}).$ (29)
We assume that a fraction $q$ of all subjects are put in treatment and $1-q$ into control. Note that as in the monopoly pricing example, this randomization can occur on either the customer or product. For simplicity, in product side randomization, we assume that for each of the $n$ types of products, a fraction $q$ of each is in treatment treatment. Let $\Pi_{0}(f_{0},f_{1},\vec{p},q)$ represent the platform’s profit from control subjects during the experiment, and $\Pi_{1}(f_{0},f_{1},\vec{p},q)$ represent the platform’s profit from treatment subjects. Let $R_{0}(f_{0},f_{1},\vec{p},q)$ be the total seller revenue for control goods, and $R_{1}(f_{0},f_{1},\vec{p},q)$ be the total seller revenue for treatment goods. We then write $\Pi_{0}(f_{0},f_{1},\vec{p},q)=f_{0}R_{0}(f_{0},f_{1},\vec{p},q)$ and $\Pi_{1}(f_{0},f_{1},\vec{p},q)=f_{1}R_{1}(f_{0},f_{1},\vec{p},q)$. We make the following assumptions for these functions, analogous to Assumption 4.2.
{assumption}
The following holds:
  1. 1.
    $R_{0}(f_{0},f_{1},\vec{p},q)$ and $R_{1}(f_{0},f_{1},\vec{p},q)$ are both differentiable over their domain.
  2. 2.
    $R_{1}(f,f,\vec{p},q)/q=R_{0}(f,f,\vec{p},q)/(1-q)=R(f,\vec{p}).$
Let $f_{1}^{n}$ be a sequence such that $\lim_{n\to\infty}f_{1}^{n}=f_{0}$. We then write the canonical estimator $\widehat{GTE}_{\Pi}$ as follows:
$\widehat{GTE}_{\Pi}(f_{0},\vec{p},q)=\lim_{n\to\infty}\frac{\Pi_{1}(f_{0},f_{1}^{n},\vec{p},q)/q-\Pi_{0}(f_{0},f_{1}^{n},\vec{p},q)/(1-q)}{f_{1}^{n}-f_{0}}.$
This is equivalent to:
$\widehat{GTE}_{\Pi}(f_{0},\vec{p},q)=\Pi_{1y}(f_{0},f_{0},\vec{p},q)/q-\Pi_{0y}(f_{0},f_{0},\vec{p},q)/(1-q),$ (30)
where $\Pi_{0y}$ and $\Pi_{1y}$ represent the derivative of the control and treatment platform profit functions with respect to the second argument respectively.
We define the bias as $Bias_{\Pi}(f_{0},\vec{p},q)=GTE_{\Pi}(f_{0},\vec{p},q)-\widehat{GTE}_{\Pi}(f_{0},\vec{p},q)$. Using Assumption 8.1 we get that $\Pi(f_{0},\vec{p},q)=\Pi_{1}(f_{0},f_{0},\vec{p},q)/q$. We can then use an analogous argument to our derivation of (11) to write bias as:
$Bias_{\Pi}(f_{0},\vec{p},q)=\Pi_{0y}(f_{0},f_{0},\vec{p},q)/(1-q)+\Pi_{1x}(f_{0},f_{0},\vec{p},q)/q.$ (31)
Using out revenue functions and equation (28) we can rewrite the preceding expressions to be more useful. First we rewrite bias as:
$Bias_{\Pi}(f_{0},\vec{p},q)=f_{0}R_{0y}(f_{0},f_{0},\vec{p},q)/(1-q)+f_{0}R_{1x}(f_{0},f_{0},\vec{p},q)/q.$
Finally, we rewrite $\widehat{GTE}_{\Pi}$ from equation (30) as follows:
$\widehat{GTE}_{\Pi}(f_{0},\vec{p},q)=R_{1}(f_{0},f_{0},\vec{p},q)/q+f_{0}R_{1y}(f_{0},f_{0},\vec{p},q)/q-f_{0}R_{0y}(f_{0},f_{0},\vec{p},q)/(1-q).$
As before, there are two necessary and sufficient conditions for the naive estimator to have the wrong sign in a fee experiment. Note that as was true in the pricing problem, bias is always positive when firms test a small price increase. This means the general conditions for a sign flip to occur are $GTE(f_{0},\vec{p})\geq 0$, meaning the true treatment effect of a small price increase is non-negative. And $Bias(f_{0},\vec{p},q)\geq\widehat{GTE}(f_{0},\vec{p},q)$, meaning there is enough bias to cause the estimated treatment effect to be non-positive.
These two conditions can be expressed as follows. First:
[…]
And second:
[…]